Circular No. 7097

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE P. Garnavich, Harvard-Smithsonian Center for Astrophysics (CfA), re-

ports for the High-Z Supernova Search Team (cf. IAUC 6160, 6646) the discovery of 13 supernovae on CCD images taken with the Cerro Tololo Interamerican Observatory (CTIO) 4-m telescope by N. B. Suntzeff, R. A. Schommer, R. C. Smith, R. Covarrubias, and A. Soderberg, CTIO; P. Challis, CfA; and B. Schmidt, Mount Stromlo and Siding Springs Observatories. The supernovae were identified by subtracting images taken in early 1998 from those obtained on Jan. 13–14 UT. The candidates were confirmed with spectra taken on Jan. 18–22 with the Keck-2 telescope by A. V. Filippenko, A. G. Riess, and D. C. Leonard (University of California, Berkeley).

SN  1999 UT  α_{2000}  δ_{2000}  Mag.  z Type
1999K Jan. 13 8^{h}03^{m}06.^{s}06 + 3^{o}24′51′′.6 21.5 R 0.04 II
1999L Jan. 13 9 28 31.96 − 0 19 35.9 23.2 R 0.65 Ia
1999M Jan. 13 10 05 46.60 − 7 27 31.3 23.4 R 0.50 Ia
1999N Jan. 13 10 04 27.93 − 7 35 51.1 23.3 R 0.54 Ia
1999O Jan. 13 10 42 05.22 − 0 29 00.2 24.1 R 0.21 II
1999P Jan. 13 10 44 08.72 − 0 49 04.5 21.8 R 0.06 Ib/c
1999Q Jan. 14 8 00 45.31 + 5 31 27.4 22.3 I 0.46 Ia
1999R Jan. 14 9 28 29.26 − 4 37 20.8 23.5 I 0.68 Ia
1999S Jan. 14 9 28 08.08 − 4 58 37.6 23.8 I 0.46 Ia
1999T Jan. 14 9 27 01.77 − 5 40 22.0 24.1 I 0.69 Ia
1999U Jan. 14 9 26 42.96 − 5 37 57.8 23.1 I 0.50 Ia
1999V Jan. 14 9 29 50.15 − 4 44 18.8 23.6 I 0.2? II?
1999W Jan. 14 4 17 46.79 − 4 50 08.4 23.4 I 0.12 II

The supernovae are within 2′′ of their host galaxies except for SN 1999K, which is 7′′.6 west and 9′′.0 south of its host; SN 1999P, which is 3′′.8 east and 8′′.4 north of its host; and SN 1999R, which is 3′′.0 west and 1′′.8 north of its host. Anyone willing to aid in the photometric follow-up may obtain finder charts by sending electronic mail to pgarnavich@cfa.harvard.edu.

COMET C/1999 A1 (TILBROOK) Visual m_{1} estimates: Jan. 14.08 UT, 10.3 (A. Hale, Cloudcroft, NM,

0.41-m reflector); 16.47, 9.8 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars); 17.50, 10.3 (Seargent); 18.54, 10.6 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); 20.54, 10.6 (Pearce); 21.96, 10.7 (J. G. de S. Aguiar, Campinas, Brazil, 0.23-m reflector).

1999 January 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green